Stars Company received a non-interest bearing promissory note of QAR 26500 for 20 months from its customer. After 5 months, Union Bank discounted the note (before maturity) when the money was worth 3.10%.

Calculate the present value of this note.

1
Expert's answer
2021-01-21T20:04:31-0500

The future value is:

"FV=PV(1+i)^n"

where "PV" is the present value, "i" is the interest per compounding period, "n" is the number of compounding periods.

Then:

"1.031PV=PV(1+i)^5"

"1.031=(1+i)^5"

"i=\sqrt[5]{1.031}-1=0.006"


We have: "FV=26500"

"PV=\frac{FV}{(1+i)^n}"


"PV=\frac{26500}{(1+0.006)^{20}}= 23511.82" QAR